OpenStudy (anonymous):

[3.05]Which is greater than seventeen twenty-fifths?

OpenStudy (skullpatrol):

$$\Huge \frac{17}{25}$$

OpenStudy (skullpatrol):

?

OpenStudy (anonymous):

20/15

OpenStudy (skullpatrol):

$$\Huge \frac{17}{25} \lt1$$

OpenStudy (skullpatrol):

$$\Huge \frac{20}{15}\gt1$$

OpenStudy (skullpatrol):